# Designer Spin-Orbit Superlattices: Symmetry-Protected Dirac Cones and Hot Spots of Spin Berry Curvature in 2D van der Waals Metamaterials

### v1

@@ Abstract
The emergence of strong relativistic spin-orbit effects in low-dimensional systems provides a rich opportunity for exploring unconventional states of matter. Here, we present a route to realize highly tunable relativistic band structures based on the lateral patterning of proximity-induced spin-orbit coupling (SOC). The concept is illustrated on a 1D patterned graphene heterostructure, where the periodic spatial modulation of SOC is shown to give rise to a rich mini-band structure with massless and massive Dirac bands carrying large Berry curvature. The envisaged systems support robust spin Hall responses driven by the quantum geometry of mini-bands, which can be tailored through metasurface fabrication methods and twisting effects. These findings suggest new pathways to 2D quantum material design and low-power spintronic applications.
@@
The intertwining of orbital and spin degrees of freedom underpin a wealth of phenomena, from the formation of topological insulators to the spin Hall effect of light [1, 2, 3, 4, 5, 6, 7]. In condensed matter systems, spin-orbit coupling (SOC) is a relativistic interaction due to the motion of electrons in the electric field of the crystal lattice, which can yield spin-dependent band structures and Berry-curvature effects that strongly influence the electrodynamics of quasiparticles [8, 9]. Because the Berry curvature flux encodes global topological invariants (such as the Chern number for quantum anomalous Hall insulators), SOC is also a key mechanism behind quantized transport in 2D and 3D topological phases of matter [10, 11].
@@
Broken symmetries alter the spin-orbital character of electronic states [12, 13], and therefore provide pathways by which to realize novel spin phenomena. Among these, the emergence of spin textures in spin-orbit-coupled systems with broken spatial inversion symmetry has generated enormous excitement in the fields of spintronics and magnonics recently [14, 15]. Owing to a close interplay of spin, lattice (pseudospin), and orbital degrees of freedom, SOC manifests both in real and momentum spaces—spin-momentum locking of spin-split Fermi surfaces [16, 17, 18], magnetic skyrmions [19, 20, 21], and persistent spin helices [22, 23, 24] are prominent examples—and forms the basis of several transport effects of fundamental and practical interest. Chief among these is the current-driven spin polarization that occurs in non-magnetic conductors with nontrivial spin textures, such as spin-momentum-locked Rashba interfaces and topological surfaces [25, 26, 27]. The ensuing net spin polarizations are often large (allowing current-induced magnetization switching of ferromagnets [28, 29, 30]) and tend to lie perpendicularly to the applied electric field owing to the tangential nature of conventional Rashba-type spin textures. Moreover, recent studies have found that the net spin orientation can be tuned in chiral materials boasting more exotic spin textures due to fully broken reflection symmetries [31, 32, 33, 34], which has the potential to unlock unconventional spin-orbit torques [35, 36, 37].
@@
Likewise, the rich landscape of spin Hall effects (SHEs) reflects the symmetries underlying spin-orbit-coupled matter [38]. Of recent and growing interest is the SHE in vertical heterostructures built from graphene and 2D semiconductors [39]. In these systems, the interfacial breaking of point-group symmetries leads to two main types of SOC that can be either induced or greatly enhanced via proximity effects: the sublattice-staggered SOC (responsible for the valley-Zeeman effect) and the more familiar Rashba SOC [40, 41, 42]. Beyond featuring an exceptionally high degree of SOC tunability via strain and twisting effects [43, 44, 45, 46, 47, 48], proximitized 2D crystals support a robust extrinsic SHE driven by scalar impurities, having no counterpart in other, non-Dirac 2D systems [49] (for a recent review see Ref. [42]). Such symmetry-breaking effects are also of ubiquitous importance for 2D quantized transport [50, 51, 52, 53], as well as for metallic anomalous Hall and magnetic spin Hall phases [54, 55].
@@
Despite this, most theoretical work so far has focused on translation invariant spin-orbit fields that reflect the periodicity of the underlying crystal structure, since this is the most conspicuous case. An interesting exception is the modulation of the strength of Rashba and Dresselhaus SOC induced in quantum-wire setups, previously explored in the context of spin-transistor devices [56, 57, 58]. Inspired by recent advances in the realization of artificial Dirac band structures in graphene with 1D superlattice potentials [59, 60, 61], the purpose of this work is to show that the quantum geometry and electrodynamic response of 2D materials can be engineered via synthetic spin-orbit fields created by a metasurface. Our proposal, outlined in Fig. 1(a), leverages proximity-induced effects between atomically thin crystals to engender effective spin-orbit fields with periodicity $a_{\text{S}}$ much greater than the lattice scale, which we call super-spin-orbit fields (SSOFs). We envision that the long-wavelength modulation of the spin-orbit field acting on charge carriers can be achieved by placing graphene on a patterned high-SOC substrate, akin to the patterning of electrostatic potentials in a lateral graphene superlattice [59, 60, 61] (other possibilities are discussed below). As we shall see, the envisaged synthetic SSOFs not only lead to the formation of mini-bands, but remarkably also impact their underlying quantum geometry, yielding a number of interesting and useful effects.
@@
$H=v\,(\bm{\sigma}\cdot\mathbf{p})\otimes s_{0}+U(x)\sigma_{0}\otimes s_{0}+H_{\text{so}}(x),$ (1)
where $v$ is the bare Fermi velocity of 2D massless Dirac fermions ($v=10^{6}$ m/s), $\sigma_{a}$ and $s_{a}$ ($a=x,y,z$) are Pauli matrices acting on the pseudospin and spin subspaces, respectively, $\sigma_{0}$ and $s_{0}$ are $2\times 2$ identity matrices, and $\mathbf{p}=-i\hbar\bm{\nabla}$ is the momentum operator. For the broad class of effective Dirac Hamiltonians that are locally invariant under the $C_{3v}$ point group [41, 42], the SSOF receives up to 3 contributions, namely, a spin-flip Rashba term [$H_{R}(x)=\lambda_{\textrm{R}}\,\Phi(x)\,(\sigma_{x}\otimes s_{y}-\sigma_{y}\otimes s_{x})$], a valley-Zeeman term due the broken sublattice symmetry [$H_{\textrm{vz}}(x)=\tau\lambda_{\textrm{vz}}\,\Phi(x)\,\sigma_{0}\otimes s_{z}$, where $\tau=\pm 1$ is the valley index], and a Kane-Mele (KM) term [$H_{\textrm{KM}}(x)=\lambda_{\textrm{KM}}\,\Phi(x)\,\sigma_{z}\otimes s_{z}$]. Here, $\lambda_{\textrm{R}}$, $\lambda_{\textrm{vz}}$ and $\lambda_{\textrm{KM}}$ denote the nominal strength of Rashba, valley-Zeeman and intrinsic-like SOC, respectively, while $\Phi(x)$ describes the spatial profile of the SOC modulation [$\Phi(x+m\,a_{\text{S}})=\Phi(x)$ with $m$ an integer]. An example of a graphene system subject to a SSOF with a square-wave profile is depicted in Fig. 1(a). Recent measurements [47, 45, 46] have shown that proximity-induced Rashba SOC in graphene/WSe2 attains giant values of up to $15$ meV [46], which is more than 350 times larger than graphene’s intrinsic SOC [63], and makes group VI dichalcogenides ideal high-SOC substrates for our proposal. Lastly, we assume that the superlattice potential, when present, is designed to track the SSOF modulation (e.g. via a patterned bottom gate), and thus write $U(x)=u\,\Phi(x)$, where $u$ is the scalar potential amplitude.
Figure 1: (a) 1D periodic modulation of the proximity-induced SOC. In this example, the SSOF is imprinted on graphene via the use of a dielectric metasurface decorated with 2D semicondutors (labelled TMD). (b) Energy dispersion of low-lying states around the $K$ point for a zero-mean square-wave profile with $a_{\text{S}}=100$ nm, $\lambda_{\text{KM}}=20$ meV, and $u=0$. (c) Same as (b) but with $u=15$ meV. As a guide to the eye, the bare energy dispersion of graphene is shown in red (inner cones).
@@
Results.—To build intuition, we first consider a SSOF with zero spatial average ($\langle\Phi(x)\rangle=0$) that locally preserves all the spatial symmetries of the honeycomb lattice, i.e. with a single term ($H_{\text{KM}}$). The energy spectrum is two-fold spin degenerate in this case and exhibits the typical mini-band structure due to a synthetic periodic perturbation. In Fig. 1(b), we show numerically exact results for a long-period square-wave modulation of type KM. The most striking feature of the low-energy spectrum is the band touching at zero energy, i.e. the SOC spatial modulation precludes the opening of a topological gap [3]. (Higher-energy mini-bands are located at energies ±$\approx\pm 2\pi v\hbar/a_{\text{S}}\approx\pm 40$ meV and thus lie outside the energy range of Fig. 1.) Importantly, the linearly dispersing zero energy states in our system cannot be gapped out without breaking the global average condition of the periodic perturbation. In other words, the Dirac point degeneracy survives SSOFs with $\langle\Phi(x)\rangle=0$. What is more, the emergent 2D Dirac fermions unveiled here remain massless even for SSOFs that locally break one or more spatial symmetries, such as a spatially modulated Rashba SOC. The robustness of the crossing point between the electron and hole mini-bands hints at interesting quantum geometry effects to be discussed shortly.
@@
Next, we observe that the KM-SSOF renormalizes the group velocity along the modulation direction, $\hat{x}$, while it produces no change perpendicularly to it. This is the opposite behavior of graphene under a periodic (scalar) potential [59], and provides a simple mechanism to fine tune charge carrier propagation. This possibility is highlighted in Fig. 1(c), showing that the combined action of a SSOF and a periodic potential squeezes the Dirac cones along both parallel and perpendicular directions to the reciprocal superlattice vector. Perturbation theory provides further insights [69]. In the limit $u,\lambda_{\text{KM}}\ll\hbar v/a_{\textrm{S}}$, the component of the group velocity parallel to the wavevector k of the low-lying Dirac states is found to be
@@
$v_{\hat{\mathbf{k}}}\cong v\left[1-\mu\,\frac{u^{2}\sin^{2}\theta_{\mathbf{k}}+\lambda_{\text{KM}}^{2}\cos^{2}\theta_{\mathbf{k}}}{\hbar^{2}v^{2}(2\pi/a_{\text{S}})^{2}}\right],$ (2)
@@
where $\theta_{\mathbf{k}}$ is the wavevector angle and $\mu$ is a geometric factor ($\mu\approx 1.645$ for a square-wave SSOF). Equation (2) shows that the periodic perturbation can be tuned to yield an isotropic group velocity. Indeed, setting $u=\pm\lambda_{\text{KM}}$ results in isotropic Dirac cones, thus mimicking the low-energy physics of bare graphene without SOC. The situation becomes richer when considering realistic systems with broken spatial symmetries as shown below. For example, Dirac fermions with isotropic behavior can be engineered by means of a pure Rashba SSOF, bypassing the need for a scalar periodic potential.
Realistic SSOFs and quantum geometry effects.—Now, we turn to the class of SSOFs that admix valley-Zeeman ($H_{\textrm{vz}}$) and Rashba ($H_{\textrm{R}}$) terms due to the breaking of spatial symmetries. Unlike the KM-type SOC in the example above, both $\lambda_{\textrm{R}}$ and $\lambda_{\textrm{vz}}$ can reach experimentally relevant energy scales, which is ideal for our proposal. We primarily focus on pure Rashba SSOFs [70], which can be realized via twist-angle engineering in graphene-on-transition metal dichalcogenides [43, 44, 45, 46]. The idea is to tune the twist angle, so that the effective SOC of charge carriers on $A$ and $B$ sublattices coincide, yielding a vanishing valley-Zeeman effect, $\lambda_{\text{vz}}=(\lambda_{A}-\lambda_{B})/2=0$. The resulting SOC is thus of Rashba type (allowed by the broken $z\rightarrow-z$ symmetry) [43, 44]. This intriguing possibility has been confirmed experimentally via quasiparticle interference imaging [46], showing that $\lambda_{\text{vz}}\approx 0$ and $\lambda_{R}\approx 15$ meV for $30^{\circ}$ twist-angle graphene-on-WSe2 systems. Armed with this important insight, we start by investigating the electronic structure induced by a square-wave Rashba SSOF. The energy dispersion of charge carriers in the three lowest-lying bands, above and below the charge neutrality point, is shown in Fig. 2(a). The calculated spectrum contains several genuine fingerprints of the SSOFs proposed in this work. Similar to the case above, the zero energy modes exhibit linear dispersion (i.e. the Dirac point degeneracy is protected). Furthermore, the behavior is isotropic. Thus, with regards to energy dispersion, this system emulates pristine graphene with a strongly renormalized Fermi velocity (see below). The massless nature of low-energy excitations is a robust feature of the 2D van der Waals metamaterials underpinning the SSOFs. In fact, only perturbations breaking the zero-average condition ($\langle\Phi\rangle=0$) can gap out the massless Dirac states [69]. As such, the zero-energy modes can be mode as robust as desired in a realistic setup, by ensuring during that the fabrication method preserves the global average of the periodic perturbation. This confers protection against local SOC fluctuations that are unavoidable in realistic systems.
@@
Figure 2: Mini-band dispersion and quantum geometry effects induced by a square-wave Rashba SSOF. (a) Energy dispersion along a cut with $k_{x}=0$. (b) Berry curvature of bands near the Fermi level along the same k-path. Mini-bands are labelled by integers next to curves [positive (negative) $n$ labels conduction (valence) bands]. Other parameters: $a_{\text{S}}=100$ nm and $\lambda_{\text{R}}=20$ meV.
@@
$\varepsilon_{{\bf k}s}^{(n=\pm 1)}\approx\pm\left(\hbar v_{\textrm{ren}}|{\bf k}|\,+s\frac{\tilde{\mu}\lambda_{\text{R}}^{3}}{\hbar^{2}v^{2}G_{1}^{4}}|{\bf k}|^{2}\right),$ (3)
where $s=\pm 1$ for spin-up (spin-down) low-energy branch, $G_{1}=2\pi/a_{\text{S}}$, $v_{\textrm{ren}}=v[1-\mu(\lambda_{\textrm{R}}/\hbar vG_{1})^{2}]$ and $\tilde{\mu}$ is a geometric factor that equals zero for sine- and square-wave modulations, but is otherwise nonzero (e.g, for Kronig-Penney-type modulations, $|\tilde{\mu}|$ attains values close to $0.26$ [69]). Hence, Rashba SSOFs with more general profiles can lift the spin degeneracy (as intuition would suggest), but only perturbatively. While the resulting spin splittings are typically small, a sizeable effect can be achieved by combining SSOFs with a periodic potential, providing a rich phenomenology for future exploration.
@@
where $v_{i}=v\,\sigma_{x}\otimes s_{0}$ and $v_{i}^{z}=v\,\sigma_{i}\otimes s_{z}$ (here, $i=x,y$) are the charge and spin velocity operators, respectively. This quantity governs the spin-Hall transport of electron wavepackets and therefore is the geometric analogue of the Berry curvature in the anomalous Hall effect [71, 55]. The SB curvature around the Dirac points is shown in Fig. 2(b). We see that the linearly dispersing zero energy modes (labelled $n=\pm 1$) are endowed with significant SB curvature, despite their massless nature. This is evidently at variance with 2D gapless Dirac systems, which have vanishing (charge) Berry curvature [10]. Thus, the emergent 2D Dirac cones reported here are not only robust against perturbations sharing the global average of the SSOFs but, remarkably, also display quantum geometric effects. Note that along the $k_{y}$-direction, the SB curvature of the massive Dirac mini-bands ($|n|=1,2$) resembles the Berry curvature of gapped 2D Dirac materials [10]. To explore this further, Figs. 3(a)-(b) show 3D plots of the SB curvature in the mini-Brillouin zone. Two features are of note: the central peaks in the SB curvature of massive Dirac mini-bands discussed earlier are identified as arising from hot spots of SB curvature [Fig. 3(b)]. On the other hand, the massless mini-bands attain a giant SB curvature at the edges of the mini-Brillouin zone ($k_{x}=\pm\pi/a_{\text{S}}$) [Fig. 3(a)], about twice as large as the Dirac-point hot spot of the massive mini-bands. We attribute this feature to the emergence of large pseudo-gaps along the SSOF direction [69]. Finally, we note that the general behavior is highly anisotropic, except in the vicinity of the Dirac point.

### v2

# Designer spin-orbit superlattices: symmetry-protected Dirac cones and spin Berry curvature in two-dimensional van der Waals metamaterials

@@ Abstract
The emergence of strong relativistic spin-orbit effects in low-dimensional systems provides a rich opportunity for exploring unconventional states of matter. Here, we present a route to realise tunable relativistic band structures based on the lateral patterning of proximity-induced spin-orbit coupling. The concept is illustrated on a patterned graphene–transition metal dichalcogenide heterostructure, where the spatially periodic spin-orbit coupling induces a rich mini-band structure featuring massless and massive Dirac bands carrying large spin Berry curvature. The envisaged systems support robust and gate-tunable spin Hall responses driven by the quantum geometry of mini-bands, which can be tailored through metasurface fabrication methods and twisting effects. These findings open pathways to two-dimensional quantum material design and low-power spintronic applications.
@@
The intertwining of orbital and spin degrees of freedom underpin a wealth of phenomena, from the formation of topological insulators to the spin Hall effect of light [1, 2, 3, 4, 5, 6, 7]. In condensed matter systems, spin-orbit coupling (SOC) is a relativistic interaction due to the motion of electrons in the electric field of the crystal lattice, which can yield spin-dependent band structures and Berry-curvature effects that strongly influence the electrodynamics of quasiparticles [8, 9]. Because the Berry curvature flux encodes global topological invariants (such as the Chern number for quantum anomalous Hall insulators), SOC is also a key mechanism behind quantised transport in topological phases of matter [10, 11].
@@
Broken symmetries alter the spin-orbital character of electronic states [12, 13], and therefore provide pathways by which to realise novel spin phenomena. Among these, the emergence of spin textures in spin-orbit-coupled systems with broken spatial inversion symmetry has generated enormous excitement in the fields of spintronics and magnonics recently [14, 15]. Owing to a close interplay of spin, lattice (pseudospin), and orbital degrees of freedom, SOC manifests both in real and momentum spaces—spin-momentum locking of spin-split Fermi surfaces [16, 17, 18], magnetic skyrmions [19, 20, 21], and persistent spin helices [22, 23, 24] are prominent examples—and forms the basis of several transport effects of fundamental and practical interest. Chief among these is the current-driven spin polarisation that occurs in non-magnetic conductors with nontrivial spin textures, such as spin-momentum-locked Rashba interfaces and topological surfaces [25, 26, 27]. The ensuing net spin polarisations are often large (allowing current-induced magnetisation switching of ferromagnets [28, 29, 30]) and tend to lie perpendicularly to the applied electric field owing to the tangential nature of conventional Rashba-type spin textures. Moreover, recent studies have found that the net spin orientation can be tuned in chiral materials boasting more exotic spin textures due to fully broken reflection symmetries [31, 32, 33, 34], which has the potential to unlock unconventional spin-orbit torques [35, 36, 37].
@@
Likewise, the rich landscape of spin Hall effects (SHEs) reflects the symmetries underlying spin-orbit-coupled matter [38]. Of recent and growing interest is the SHE in vertical heterostructures built from graphene and two-dimensional (2D) semiconductors [39]. In these systems, the interfacial breaking of point-group symmetries leads to two main types of SOC that can be either induced or greatly enhanced via proximity effects: the sublattice-staggered SOC (underlying the valley-Zeeman effect) and the more familiar Rashba SOC [40, 41, 42]. Beyond featuring an exceptionally high degree of SOC tunability via strain and twisting effects [43, 44, 45, 46, 47, 48], proximitized 2D crystals support robust extrinsic SHEs due to scalar impurities, having no counterpart in other, non-Dirac 2D systems [49, 50] (for a recent review see Ref. [42]). Such symmetry-breaking effects are also of ubiquitous importance for 2D quantised transport [51, 52, 53, 54], as well as for metallic anomalous Hall and magnetic spin Hall phases [55, 56].
@@
Despite this, most theoretical work so far has focused on translation invariant spin-orbit fields that reflect the periodicity of the underlying crystal structure, since this is the most conspicuous case. An interesting exception is the modulation of the strength of Rashba and Dresselhaus SOC induced in quantum-wire setups, previously explored in the context of spin-transistor devices [57, 58, 59]. Inspired by recent advances in the realisation of artificial Dirac band structures in graphene with one-dimensional (1D) superlattice potentials [60, 61, 62], the purpose of this work is to show that the quantum geometry and electrodynamic response of 2D materials can be engineered via synthetic spin-orbit fields created by a metasurface. Our proposal, outlined in Fig. 1(a), leverages proximity-induced effects between atomically thin crystals to engender effective spin-orbit fields with periodicity $a_{\text{S}}$ much greater than the lattice scale, which we call super-spin-orbit fields (SSOFs). We envision that the long-wavelength modulation of the spin-orbit field acting on charge carriers can be achieved by placing graphene on a patterned high-SOC substrate, akin to the patterning of electrostatic potentials in a lateral graphene superlattice [60, 61, 62] (other possibilities are discussed below). As we shall see, the envisaged synthetic SSOFs not only lead to the formation of mini-bands but also impact their underlying quantum geometry, yielding a number of useful effects. This includes the counterintuitive and exotic possibility of creating linearly dispersing spin-degenerate electronic states even for Rashba-type SSOFs where spatial inversion symmetry is strictly broken. Our proposal is, therefore, complementary to previous superlattice setups, where the presence of spatially uniform SOC components generally leads to spin-split energy bands with non-linear dispersion as well as energy gaps [63, 64, 65]. The SSOFs are also distinct from superlattices arising from the periodic modulation of on-site staggered potentials [66, 67], which lack SOC effects. Another advantage of the SSOFs introduced here is that they intrinsically generate semimetallic phases without the need for periodic Zeeman fields [68]. Moreover, similar to spatially uniform SOC, the SSOFs endow the electronic states of graphene with spin Berry curvature, thus paving the way to SHEs with unique geometric features.
@@
$H_{\tau}=v\,(\bm{\sigma}\cdot\mathbf{p})\otimes s_{0}+U(x)\sigma_{0}\otimes s_{0}+H_{\text{so},\tau}(x),$ (1)
where $v$ is the bare Fermi velocity of 2D massless Dirac fermions ($v=10^{6}$ m/s), $\sigma_{a}$ and $s_{a}$ ($a=x,y,z$) are Pauli matrices acting on the pseudospin and spin subspaces, respectively, $\sigma_{0}$ and $s_{0}$ are $2\times 2$ identity matrices, $\mathbf{p}=-i\hbar\bm{\nabla}$ is the momentum operator, and $\tau=\pm 1$ is the valley index. For the broad class of Dirac Hamiltonians that are locally invariant under the $C_{3v}$ point group [41, 42], the SSOF term $H_{\text{so},\tau}(x)$ receives up to 3 contributions, namely, a spin-flip Rashba term [$H_{R}(x)=\lambda_{\textrm{R}}\,\Phi(x)\,(\sigma_{x}\otimes s_{y}-\sigma_{y}\otimes s_{x})$], a valley-Zeeman term due the broken sublattice symmetry [$H_{\textrm{vz},\tau}(x)=\tau\lambda_{\textrm{vz}}\,\Phi(x)\,\sigma_{0}\otimes s_{z}$], and a Kane-Mele (KM) term [$H_{\textrm{KM}}(x)=\lambda_{\textrm{KM}}\,\Phi(x)\,\sigma_{z}\otimes s_{z}$]. Here, $\lambda_{\textrm{R}}$, $\lambda_{\textrm{vz}}$ and $\lambda_{\textrm{KM}}$ respectively denote the nominal strength of the Rashba, valley-Zeeman and intrinsic-like SOC induced by the application of the SSOF, while $\Phi(x)$ describes the spatial profile of the SOC modulation [$\Phi(x+m\,a_{\text{S}})=\Phi(x)$ with $m$ an integer]. We note that the presence of the periodic modulation yields a mini-band energy spectrum, ${\varepsilon_{n\mathbf{k}}}$, where $\0n\in\mathbb{Z}\backslash\{0\}$ is the mini-band index and $\mathbf{k}$ is the Bloch wavevector relative to a Dirac valley; see Supplementary Material (SM) appended below for additional details.
Figure 1: (a) 1D periodic modulation of the proximity-induced SOC. In this example, the SSOF is imprinted on graphene via the use of a dielectric metasurface decorated with 2D semiconductors (labelled TMD). (b) Energy dispersion of low-lying states around the $K$ valley for a zero-mean square-wave profile with $a_{\text{S}}=100$ nm, $\lambda_{\text{KM}}=20$ meV, and $u=0$. (c) Same as (b) but with $u=15$ meV. As a guide to the eye, the bare energy dispersion of graphene is shown in red (inner cones).
@@
Results.—To build intuition, we first consider an SSOF with zero spatial average ($\langle\Phi(x)\rangle=0$) that locally preserves all the spatial symmetries of the honeycomb lattice, i.e. with a single term ($H_{\text{KM}}$). The energy spectrum is two-fold spin degenerate in this case and exhibits the typical mini-band structure due to a synthetic periodic perturbation. In Fig. 1(b), we show numerically exact results for a long-period square-wave modulation of type KM; details are provided in the Methods. The most striking feature of the low-energy spectrum is the band touching at zero energy, i.e. the SOC spatial modulation precludes the opening of a topological gap [3]. (Higher-energy mini-bands are located at energies ±$\approx\pm 2\pi v\hbar/a_{\text{S}}\approx\pm 40$ meV and thus lie outside the energy range of Fig. 1.) Importantly, the linearly dispersing zero energy states in our system cannot be gapped out without breaking the global average condition of the periodic perturbation. In other words, the Dirac point degeneracy survives SSOFs with $\langle\Phi(x)\rangle=0$ (the reader is referred to SM for numerical and analytical evidence supporting the generality of this statement). What is more, the emergent 2D Dirac fermions unveiled here remain massless even for SSOFs that locally break one or more spatial symmetries, such as a spatially modulated Rashba SOC. The robustness of the crossing point between the electron and hole mini-bands hints at significant effects of quantum geometry, which will be discussed shortly.
@@
Next, we observe that the KM-SSOF renormalises the group velocity along the modulation direction, $\hat{x}$, while it produces no change perpendicularly to it. This is the opposite behaviour of graphene under a periodic (scalar) potential [60], and provides a simple mechanism to fine tune charge carrier propagation. This possibility is highlighted in Fig. 1(c), showing that the combined action of a SSOF and a periodic potential squeezes the Dirac cones along both parallel and perpendicular directions to the reciprocal superlattice vector. Perturbation theory provides further insights, as detailed in the SM. In the limit $u,\lambda_{\text{KM}}\ll\hbar v/a_{\textrm{S}}$, the component of the group velocity parallel to the wavevector k of the low-lying Dirac states reads as
@@
$v_{\hat{\mathbf{k}}}\cong v\left[1-\xi\,\frac{u^{2}\sin^{2}\theta_{\mathbf{k}}+\lambda_{\text{KM}}^{2}\cos^{2}\theta_{\mathbf{k}}}{\hbar^{2}v^{2}{G_{1}}^{2}}\right],$ (2)
@@
where $\theta_{\mathbf{k}}$ is the wavevector angle, $\xi$ is a geometric factor ($\xi\approx 1.645$ for a square-wave SSOF), and $G_{1}=2\pi/a_{\text{S}}$. Equation (2) shows that the periodic perturbation can be tuned to yield an isotropic group velocity. Indeed, setting $u=\pm\lambda_{\text{KM}}$ results in isotropic Dirac cones, thus mimicking the low-energy physics of bare graphene without SOC. The situation becomes richer when considering realistic systems with broken spatial symmetries as shown below. For example, Dirac fermions with isotropic behaviour can be engineered by means of a pure Rashba SSOF, bypassing the need for a scalar periodic potential.
@@
Figure 2: (a) Dispersion of low-lying mini-bands along a cut with $k_{x}=0$. (b) Spin Berry curvature along the same k-path. Mini-bands are labelled by integers next to curves [positive (negative) $n$ labels conduction (valence) bands]. Other parameters: $a_{\text{S}}=100$ nm and $\lambda_{\text{R}}=20$ meV.
Figure 3: (a) Momentum-space distribution of the SB curvature of the massless Dirac mini-bands ($n=\pm 1$). (b) Same as in (a) for the mini-bands $n=\pm 2$. SSOF parameters as in Fig. 2. (c) Spin Hall conductivity $\sigma^{z}_{yx}$ as a function of the chemical potential for $k_{B}T=26$ meV and selected twisting-induced SOC modulations, namely: pure Rashba SSOF (solid lines), admixed Rashba-valley-Zeeman SSOFs with $\lambda_{\text{vz}}=\lambda_{\text{R}}/2$ (dashed lines) and $\lambda_{\text{vz}}=\lambda_{\text{R}}$ (dot-dashed lines). Here, green (blue) curves correspond to $\lambda_{R}=10$ meV ($20$ meV).
Realistic SSOFs and quantum geometry effects.—Now, we turn to the class of SSOFs that admix valley-Zeeman ($H_{\textrm{vz}}$) and Rashba ($H_{\textrm{R}}$) terms due to the breaking of spatial symmetries. Unlike the KM-type SOC in the example above, both $\lambda_{\textrm{R}}$ and $\lambda_{\textrm{vz}}$ can reach experimentally relevant energy scales, which is ideal for our proposal. We primarily focus on pure Rashba SSOFs which can be realised via twist-angle engineering in graphene-on-transition metal dichalcogenide (TMD) heterostructures [43, 44, 45, 46]. We neglect the KM-type SOC, which due to its smallness [70] is unimportant. The idea is to tune the twist angle, so that the effective SOC of charge carriers on $A$ and $B$ sublattices coincide, yielding a vanishing valley-Zeeman effect, $\lambda_{\text{vz}}=(\lambda_{A}-\lambda_{B})/2=0$. The resulting SOC is thus of Rashba type (allowed by the broken $z\rightarrow-z$ symmetry) [43, 44]. This intriguing possibility has been confirmed experimentally via quasiparticle interference imaging [46], showing that $\lambda_{\text{vz}}\approx 0$ and $\lambda_{R}\approx 15$ meV for $30^{\circ}$ twist-angle graphene-on-WSe2 systems. Armed with this important insight, we start by investigating the electronic structure induced by a square-wave Rashba SSOF. The energy dispersion of charge carriers in the three lowest-lying bands, above and below the charge neutrality point, is shown in Fig. 2(a). The calculated spectrum contains several genuine fingerprints of the SSOFs proposed in this work. Similar to the case above, the zero energy modes exhibit linear dispersion (i.e. the Dirac point degeneracy is protected). Furthermore, the behaviour is isotropic. Thus, with regards to energy dispersion, this system emulates pristine graphene with a strongly renormalised Fermi velocity (see below). The massless nature of low-energy excitations is a robust feature of the 2D van der Waals metamaterials underpinning the SSOFs. In fact, only perturbations breaking the zero-average condition ($\langle\Phi\rangle=0$) can gap out the massless Dirac states (see SM). As such, the zero-energy modes can be mode as robust as desired in a realistic setup, by ensuring that the fabrication method preserves the global average of the periodic perturbation. This confers protection against local SOC fluctuations that are unavoidable in realistic systems.
@@
$\varepsilon_{n=\pm 1,{\bf k}s}^{(3)}\approx\pm\left(\hbar v_{\textrm{ren}}|{\bf k}|\,+s\tilde{\xi}\frac{\lambda_{\text{R}}^{3}}{\hbar^{2}v^{2}G_{1}^{4}}|{\bf k}|^{2}\right),$ (3)
where $s=\pm 1$ for spin-up (spin-down) low-energy branch, $v_{\textrm{ren}}=v[1-\xi(\lambda_{\textrm{R}}/\hbar vG_{1})^{2}]$ and $\tilde{\xi}$ is a geometric factor that equals zero for sine- and square-wave modulations, but is otherwise non-zero (e.g, for Kronig-Penney-type modulations, $|\tilde{\xi}|$ attains values close to $0.26$; see SM for additional details. Hence, Rashba SSOFs with more general profiles can lift the spin degeneracy (as intuition would suggest), but only perturbatively. While the resulting spin splittings are typically small, a sizeable effect can be achieved by combining SSOFs with a periodic potential, providing a rich phenomenology for future exploration.
@@
where $v_{i}=v\,\sigma_{i}\otimes s_{0}$ and $v_{i}^{z}=v\,\sigma_{i}\otimes s_{z}$ (here, $i=x,y$) are the charge and spin velocity operators, respectively. This quantity governs the spin Hall transport of electron wavepackets and therefore is the geometric analogue of the Berry curvature in the anomalous Hall effect [76, 56]. The SB curvature around the Dirac points is shown in Fig. 2(b). We see that the linearly dispersing zero energy modes (labelled $n=\pm 1$) are endowed with significant SB curvature, despite their massless nature. This is evidently at variance with 2D gapless Dirac systems, which have vanishing (charge) Berry curvature [10]. Thus, the emergent 2D Dirac cones reported here are not only robust against perturbations sharing the global average of the SSOFs but also display quantum geometric effects. To explore this further, Figs. 3(a)-(b) show 3D plots of the SB curvature of mini-bands with $n=\pm 1,\pm 2$ in the mini-Brillouin zone. Two features are of note. First, the central peaks in the SB curvature of the massive Dirac mini-bands $n=\pm 2,\pm 3$ discussed earlier are seen to arise from local hot spots of SB curvature [see the local maximum of $\Omega_{yx,\pm 2}^{z}$ at $\textbf{k}=0$ in Fig. 3(b)]. The SB curvature also displays pronounced local minima near the zone edges, where $\Omega_{yx,\pm 2}$ attains large negative values. Second, the massless mini-bands ($n=\pm 1$) have a giant SB curvature at the edges of the mini-Brillouin zone ($k_{x}=\pm\pi/a_{\text{S}}$) [see Fig. 3(a)], about twice as large as the Dirac-point hot spot of the massive mini-bands. We attribute this curious feature to the emergence of large pseudo-gaps along the SSOF direction; see SM. Finally, we note that the general behaviour is highly anisotropic, except in the immediate vicinity of the Dirac point.